# Line bundles and exact sequences for the ideal class group and the Picard group

### v1

@@ Abstract.
For any extension of commutative rings $A\subseteq B$ we first naturally define a group $\operatorname{\mathfrak{C}}(A,B)$, that we call the ideal class group of this extension (we will see that both the classical ideal class group and, surprisingly, the Picard group are special cases of this structure), then, as a first main result, we obtain the following exact sequence of Abelian groups:
In the second main theorem, using ideas from algebraic geometry, we prove a general result which asserts that every commutative ring has a ring extension whose Picard group is trivial. These two theorems have several interesting applications. For instance, we show that the Picard group of every commutative ring is canonically isomorphic to the ideal class group of a certain extension of that ring. As another application, we prove that every line bundle has the avoidance property. The next main result asserts that, for any extensions of commutative rings $A\subseteq B\subseteq C$, we have the following exact sequence of Abelian groups:
We show that for any extension of commutative rings $A\subseteq B$, we have a canonical isomorphism of groups $\operatorname{\mathfrak{C}}(A,B)\simeq\operatorname{\mathfrak{C}}(A_{\mathrm{red}},B_{\mathrm{red}})$. We also prove that if an extension of rings has a retraction, then its ideal class group is trivial. In the next key result, for any ring $A$, we obtain a canonical group map $\mathfrak{C}(A)\rightarrow\mathfrak{C}(A_{\mathrm{red}})$ which is always injective. In the same vein, we find an important example showing that this canonical injection is not necessarily surjective, and hence in general the ideal class group does not behave like the Picard group.
@@ 1. Introduction
The classical ideal class group of an integral domain is of special importance in algebraic number theory and understanding this group is indeed one of the central goals of this field. The first main goal of this article is to generalize naturally this classical group to the more general case of an arbitrary extension of commutative rings. In fact, for any extension of commutative rings $A\subseteq B$, just as in the classical construction, we define $\operatorname{\mathfrak{C}}(A,B)$ as the group of invertible ideals of this extension modulo the subgroup of principal invertible ideals. We call $\operatorname{\mathfrak{C}}(A,B)$ the ideal class group of the extension $A\subseteq B$.
@@
Next, in Theorem 4.2, we discovered an unexpected result which asserts that every commutative ring has a ring extension whose Picard group is trivial. This result had an exciting story for us. Indeed, at first we formulated this theorem merely as a conjecture. We then worked on this problem for a while. Later, we (the author and Melvin Hochster) found an example of a commutative ring that cannot be extended to a semi-local ring (i.e., a ring with finitely many maximal ideals). We thought that this ring would probably be a counterexample to our conjecture, considering that every semi-local ring has a trivial Picard group. Then, relying on this example, we were almost convinced that our conjecture was most likely not true.
But to be completely sure, we then asked Pierre Deligne whether this conjecture was true or not. But later Deligne, quite surprisingly, answered that the conjecture was correct and sent us a geometrical proof of it. However, his very sketchy proof was too difficult for us to understand and was based entirely on modern algebraic geometry. In fact, his proof was based on the following two main ideas. The first idea was that line bundles (projective modules everywhere of rank one) are classified by $\mathbb{G}_{\mathrm{m}}$-torsors (principal homogeneous spaces). The second idea was that the universal way to vanish (trivialize) a torsor under a group scheme is to base change by this same torsor.
Then, using the algebraic interpretations of these geometric ideas, we (the author and Hochster) were able to transform Deligne’s proof into a purely algebraic proof. In fact, from the algebraic point of view, one of the key ideas of the proof is that the push-out of every line bundle is vanished (trivialized) by its own torsor algebra. For any line bundle $L\in\operatorname{Pic}(A)$, we call the $\mathbb{Z}$-graded commutative ring $\bigoplus\limits_{n\in\mathbb{Z}}L^{\otimes n}$ the torsor algebra of $L$ (this algebra plays an important role in our proof). The second main idea is that, by taking the tensor products of the torsor algebras vanishing each of the line bundles, one can vanish all them together (then repeat this process to vanish the new line bundles which could have appeared). The final idea of the proof is that the Picard group commutes with direct limits.
In Theorem 4.4, we prove another important result which asserts that for a given extension of rings $A\subseteq B$, if $A$ has finitely many maximal ideals then every invertible ideal of this extension is principal. Then some applications of the above theorems are given.
As a first and major application, we show that the Picard group of every ring with finitely many maximal ideals is trivial. We also show that in a given extension of rings $A\subseteq B$, if $B$ has finitely many maximal ideals then we have the canonical isomorphism of groups $\operatorname{\mathfrak{C}}(A,B)\simeq\operatorname{Pic}(A)$. As another application, for any commutative ring $A$, we obtain the following exact sequence of Abelian groups:
@@
where $\operatorname{\mathfrak{C}}(A)=\operatorname{\mathfrak{C}}(A,T(A))$ and $T(A)$ is the total ring of fractions of $A$. In particular, if $A$ is a reduced ring with finitely many minimal primes, then we have the canonical isomorphism of groups $\operatorname{\mathfrak{C}}(A)\simeq\operatorname{Pic}(A)$.
These applications culminate in showing that the Picard group is a special case of the ideal class group. More precisely, we show that the Picard group of every commutative ring is canonically isomorphic to the ideal class group of a certain extension of that ring (see Corollary 4.3). This result also highlights the importance of our new concept of the ideal class group of an extension. As another important application, in Corollary 4.10, we prove that every line bundle has the avoidance property. This result considerably generalizes [7, Theorem 1.5] and our recent result [2, Corollary 3.2].
@@
Next we will examine the vanishing aspect of the ideal class group. In this regard, in Theorem 5.10, we prove a general result which asserts that if an extension of rings has a retraction, then its ideal class group is trivial. As an application, we show that for any commutative ring $A$ and for any commutative monoid $M$, the ideal class group of the monoid-ring extension $A\subseteq A[M]$ is trivial. In particular, the ideal class groups of the polynomial ring extensions $A\subseteq A[x]$ and $A\subseteq A[x,x^{-1}]$ are trivial. The ideal class group of the extension $A\subseteq A[[x]]$ is also trivial. In other words, every invertible ideal of these extensions is principal.
Finally, we give another important example showing that the kernel of the canonical group map $\operatorname{Pic}(A)\rightarrow\operatorname{Pic}(B)$, induced by a ring extension $A\subseteq B$, is very sensitive to changes in scalars (see Example 5.13).
@@ 2. Preliminaries
In this section, we recall some basic background for the convenience of the reader. The subject of this article can be seen at the intersection of commutative algebra, algebraic geometry, algebraic K-theory, and algebraic number theory. The group of units (invertible elements) of a commutative ring $A$ is denoted by $A^{\ast}$. The set of zero-divisors of $A$ is denoted by $Z(A)$. By $A_{\mathrm{red}}$ we mean the ring $A$ modulo its nil-radical.
@@
If $M$ is a finitely generated flat module over a ring $A$, then for each $\mathfrak{p}\in\operatorname{Spec}(A)$, there exists a (unique) natural number $n_{\mathfrak{p}}\geqslant 0$ such that $M_{\mathfrak{p}}\simeq(A_{\mathfrak{p}})^{n_{\mathfrak{p}}}$ as $A_{\mathfrak{p}}$-modules, because it is well known that every finitely generated flat module over a local ring is a free module (see [6, Theorem 7.10]). In fact, this number $n_{\mathfrak{p}}$ is the dimension of $\kappa(\mathfrak{p})$-vector space $M\otimes_{A}\kappa(\mathfrak{p})$ where $\kappa(\mathfrak{p})=A_{\mathfrak{p}}/\mathfrak{p}A_{\mathfrak{p}}$ is the residue field of $A$ at $\mathfrak{p}$. Hence, we obtain a function $\mathrm{r}_{M}:\operatorname{Spec}(A)\rightarrow\mathbb{Z}$ given by $\mathfrak{p}\mapsto\operatorname{rank}_{R_{\mathfrak{p}}}(M_{\mathfrak{p}})=n_{\mathfrak{p}}$. This function is called the rank map of $M$. It is well known that the rank map of a finitely generated flat $A$-module is continuous if and only if it is a projective $A$-module.
Let $A$ be a ring. It is well known that an $A$-module $M$ is a finitely generated projective $A$-module of constant rank 1 if and only if there are (finitely many) elements $f_{1},\ldots,f_{n}\in A$ which generate the unit ideal $A$ and $M_{f}$ is a free $A_{f}$-module of constant rank 1 for all $f\in\{f_{1},\ldots,f_{n}\}$, or equivalently, the canonical morphism of $A$-modules $M\otimes_{A}\widehat{M}\rightarrow A$ given by $x\otimes f\mapsto f(x)$ is an isomorphism where $\widehat{M}=\operatorname{Hom}_{A}(M,A)$ is the dual of $M$. In this case, $\widehat{M}$ is also a finitely generated projective $A$-module of rank 1. All these results can be found in standard books on algebraic K-theory (see e.g. [1], [10], [13] and [16]).
For any commutative ring $A$, by $\operatorname{Pic}(A)$ we mean the set of isomorphism classes of line bundles over $A$ (finitely generated projective $A$-modules of constant rank 1). This set $\operatorname{Pic}(A)$ by the operation $L_{1}]\cdot[L_{2}]=[L_{1}\otimes_{A}L_{2}]$ is an Abelian group whose identity element is the isomorphism class of $A$ and the inverse of each element $L]\in\operatorname{Pic}(A)$ is the isomorphism class of $\widehat{L}$. The group $\operatorname{Pic}(A)$ is called the Picard group of $A$. In this group $\operatorname{Pic}(A)$, we will often denote the isomorphism class $[L]$ simply by $L$ if there is no confusion.
@@
It can be seen that the Picard group commutes with direct limits. More precisely, if $(A_{k},f_{i,k})$ is a direct system of rings over a (directed) poset $I$, then the groups $\operatorname{Pic}(A_{k})$ together with the maps $\operatorname{Pic}(f_{i,k})$, as transition morphisms, is a direct system of Abelian groups over the same (directed) poset and we have the canonical isomorphism of groups $\operatorname{Pic}(\lim\limits_{\overrightarrow{k\in I}}A_{k})\simeq\lim\limits_{\overrightarrow{k\in I}}\operatorname{Pic}(A_{k})$. For given rings $A$ and $B$ consider their direct product ring $R=A\times B$, then the map $\operatorname{Pic}(R)\rightarrow\operatorname{Pic}(A)\times\operatorname{Pic}(B)$ given by $L\mapsto(L\otimes_{R}A,L\otimes_{R}B)$ is an isomorphism of groups.
@@
Let $I$ be an ideal of a ring $A$ and consider the canonical ring map $f:A\rightarrow A/I$. If $I$ is contained in the Jacobson radical of $A$, then the map $\operatorname{Pic}(f):\operatorname{Pic}(A)\rightarrow\operatorname{Pic}(A/I)$ is injective. If moreover, $A$ is complete with respect to the $I$-adic topology or $I$ is contained in the nil-radical of $A$, then the map $\operatorname{Pic}(f)$ is an isomorphism.
Although we could not find a reference that provides an integrated and single proof of this theorem, its proof can be found in the combination of [1, Chap III, Proposition 2.12], [5, Chap X, §5, Theorem 5.10, 1.(b)] and [16, Chap I, §2, Ex.2.2].
In this article we will need only a special but important case of the above theorem which asserts that: for any ring $A$ then the group morphism $\operatorname{Pic}(A)\rightarrow\operatorname{Pic}(A_{\mathrm{red}})$, induced by the canonical ring map $A\rightarrow A_{\mathrm{red}}$, is an isomorphism.

### v3

# Line bundles, ideal class group of an extension and Picard group

@@ Abstract.
For any extension of commutative rings $A\subseteq B$ we first naturally define a group $\operatorname{\mathfrak{C}}(A,B)$, that we call the ideal class group of this extension. Then we study the basic properties of this group.
Next, using ideas from algebraic geometry, we prove that every commutative ring has a (faithfully flat) ring extension whose Picard group is trivial.
These results have several interesting applications. In particular, we will see that both the classical ideal class group and, surprisingly, the Picard group are special cases of this new structure of the ideal class group of an extension. As another application, we show that every line bundle has the avoidance property.
The Picard group does not necessarily commute with the “infinite” direct product of rings. However, we show that the Picard group of the direct product of an (infinite) family of rings can be canonically embedded in the direct product of the Picard groups of the corresponding factor rings. We also prove that the Picard group of every zero-dimensional ring is trivial.
Among the next main results, we prove that for any extension of commutative rings $A\subseteq B$, we have a canonical isomorphism of groups $\operatorname{\mathfrak{C}}(A,B)\simeq\operatorname{\mathfrak{C}}(A_{\mathrm{red}},B_{\mathrm{red}})$.
Next, we prove that the ideal class group of any extension of rings $A\subseteq B$ with $A$ has finitely many maximal ideals is trivial as well.
In the next key result, for any ring $A$, we obtain a canonical group map $\mathfrak{C}(A)\rightarrow\mathfrak{C}(A_{\mathrm{red}})$ which is always injective. In the same vein, we find an important example showing that this canonical injection is not necessarily surjective, and hence in general the ideal class group does not behave like the Picard group.
@@ 1. Introduction
The subject of this article lies at the fascinating intersection of commutative algebra, algebraic geometry, and algebraic number theory.
The classical ideal class group of an integral domain is of special importance in algebraic number theory and understanding the structure of this group is indeed one of the central goals of this field. The first main aim of this article is to generalize naturally this classical group to the more general setting of an arbitrary extension of commutative rings. In fact, for any extension of commutative rings $A\subseteq B$, just as in the classical construction, we define $\operatorname{\mathfrak{C}}(A,B)$ as the group of invertible ideals of this extension modulo the subgroup of principal invertible ideals. We call $\operatorname{\mathfrak{C}}(A,B)$ the ideal class group of the extension $A\subseteq B$.
@@
Next, in Theorem 4.2, we obtained an unexpected result which asserts that every commutative ring has a faithfully flat ring extension whose Picard group is trivial.
This result had an exciting story for us. Indeed, at first we formulated this theorem merely as a conjecture. Next, we asked Pierre Deligne whether this conjecture was true or not. Then, he answered that the conjecture was correct and sent us a geometrical proof of it. However, his very sketchy proof was too difficult for us to understand and was based entirely on modern algebraic geometry. In fact, his proof was based on the following two main ideas. The first idea was that line bundles (projective modules everywhere of rank one) are classified by $\mathbb{G}_{\mathrm{m}}$-torsors (principal homogeneous spaces). The second idea was that the universal way to vanish (trivialize) a torsor under a group scheme is to base change by this same torsor.
Then, using the algebraic interpretations of these geometric ideas, we (with the collaboration of Melvin Hochster) were able to transform Deligne’s proof into a purely algebraic proof. In fact, from the algebraic point of view, one of the key ideas of the proof is that the push-out of every line bundle is vanished (trivialized) by its own torsor algebra. For any line bundle $L\in\operatorname{Pic}(A)$, we call the $\mathbb{Z}$-graded commutative ring $\bigoplus\limits_{n\in\mathbb{Z}}L^{\otimes n}$ the torsor algebra of $L$ (this algebra plays an important role in our proof). The second main idea is that, by taking the tensor products of the torsor algebras vanishing each of the line bundles, one can vanish all them together (then repeat this process to vanish the new line bundles which could have appeared). The final idea of the proof is that the Picard group commutes with direct limits.
Next, in Theorem 4.6, we prove that the Picard group of the direct product of an (infinite) family of rings can be canonically embedded in the direct product of the Picard groups of the factor rings. In particular, the Picard group of the direct product of an arbitrary family of rings is trivial if and only if the Picard group of each factor ring is trivial. As an application, a weak version of the above theorem is easily deduced which states that every commutative ring has a ring extension whose Picard group is trivial.
In Theorem 4.4, we prove another important result which asserts that for a given extension of rings $A\subseteq B$, if $A$ has finitely many maximal ideals then every invertible ideal of this extension is principal.
Then some applications of the above theorems are given. We show that in a given extension of rings $A\subseteq B$, if $B$ has finitely many maximal ideals then we have a canonical isomorphism of groups $\operatorname{\mathfrak{C}}(A,B)\simeq\operatorname{Pic}(A)$. As another application, for any commutative ring $A$, we obtain the following exact sequence of Abelian groups:
@@
where $\operatorname{\mathfrak{C}}(A):=\operatorname{\mathfrak{C}}(A,T(A))$ and $T(A)$ is the total ring of fractions of $A$. In particular, if $A$ is a reduced ring with finitely many minimal primes (e.g. an integral domain), then we have a canonical isomorphism of groups $\operatorname{\mathfrak{C}}(A)\simeq\operatorname{Pic}(A)$.
These applications culminate in showing that the Picard group is a special case of the ideal class group. More precisely, we show that the Picard group of every commutative ring is canonically isomorphic to the ideal class group of a certain extension of that ring (see Corollary 4.3). These applications indeed show that this new concept of the ideal class group of an extension unifies classical structures.
As another important application, in Corollary 4.14, we prove that every line bundle has the avoidance property. This result considerably generalizes [7, Theorem 1.5] and our recent result [2, Corollary 3.2].
In Theorem 4.17, we prove another interesting result which asserts that the Picard group of every zero-dimensional ring is trivial.
@@
Next we will examine the vanishing aspect of the ideal class group. In this regard, in Corollary 5.10, we show that if an extension of rings has a retraction, then its ideal class group is trivial. As an application, we show that for any commutative ring $A$ and for any commutative monoid $M$, the ideal class group of the monoid-ring extension $A\subseteq A[M]$ is trivial. In particular, the ideal class groups of the polynomial ring extensions $A\subseteq A[x]$ and $A\subseteq A[x,x^{-1}]$ are trivial. The ideal class group of the extension $A\subseteq A[[x]]$ is also trivial. In this regard, we also show that the ideal class group of any extension of rings $A\subseteq B$ with $A$ has finitely many maximal ideals is trivial as well (see Theorem 4.4). In other words, every invertible ideal of these extensions is principal.
Finally, we give another important example showing that there are some ring extensions $A\subseteq B$ such that the kernel of the canonical group map $\operatorname{Pic}(A)\rightarrow\operatorname{Pic}(B)$ is very sensitive to changes in scalars (see Example 5.12).
@@ 2. Preliminaries
In this section, we recall some basic background for the convenience of the reader. The group of units (invertible elements) of a commutative ring $A$ is denoted by $A^{\ast}$. The set of zero-divisors of $A$ is denoted by $Z(A)$. By $A_{\mathrm{red}}$ we mean the ring $A$ modulo its nil-radical. If $r$ is an element of a ring $A$, then $D(r)=\{\mathfrak{p}\in\operatorname{Spec}(A):r\notin\mathfrak{p}\}$.
@@
If $M$ is a finitely generated flat module over a ring $A$, then for each $\mathfrak{p}\in\operatorname{Spec}(A)$, there exists a (unique) natural number $n_{\mathfrak{p}}\geqslant 0$ such that $M_{\mathfrak{p}}\simeq(A_{\mathfrak{p}})^{n_{\mathfrak{p}}}$ as $A_{\mathfrak{p}}$-modules (because it is well known that every finitely generated flat module over a local ring is a free module, see e.g. [6, Theorem 7.10]). Hence, we obtain a function $\mathrm{r}_{M}:\operatorname{Spec}(A)\rightarrow\mathbb{Z}$ given by $\mathfrak{p}\mapsto n_{\mathfrak{p}}$. We call $\mathrm{r}_{M}$ the rank map of $M$.
It can be also seen that this number $\operatorname{rank}_{R_{\mathfrak{p}}}(M_{\mathfrak{p}}):=n_{\mathfrak{p}}$ is the dimension of $\kappa(\mathfrak{p})$-vector space $M\otimes_{A}\kappa(\mathfrak{p})$ where $\kappa(\mathfrak{p})=A_{\mathfrak{p}}/\mathfrak{p}A_{\mathfrak{p}}$ is the residue field of $A$ at $\mathfrak{p}$.
If $M$ and $N$ are finitely generated flat $R$-modules, then $M\oplus N$ and $M\otimes_{R}N$ are finitely generated flat $R$-modules and we have $\mathrm{r}_{M\oplus N}=\mathrm{r}_{M}+\mathrm{r}_{N}$ and $\mathrm{r}_{M\otimes_{R}N}=\mathrm{r}_{M}\cdot\mathrm{r}_{N}$.
It is well known that the rank map of a finitely generated flat $A$-module is continuous if and only if it is a projective $A$-module.
By invertible module over a ring $A$ we mean a finitely generated projective $A$-module of constant rank 1.
Let $M$ and $N$ be modules over a ring $A$. If $M\otimes_{A}N\simeq A^{n}$ as $A$-modules for some $n\geqslant 1$, then $M$ and $N$ are finitely generated projective $A$-modules. In particular, if $n=1$, then $M$ and $N$ are invertible $A$-modules.
It can be seen that a module $M$ over a ring $A$ is invertible (i.e., finitely generated projective $A$-module of constant rank 1) if and only if there are (finitely many) elements $f_{1},\ldots,f_{n}\in A$ which generate the unit ideal of $A$ and $M_{f}$ is a free $A_{f}$-module of constant rank 1 for all $f\in\{f_{1},\ldots,f_{n}\}$, or equivalently, the canonical morphism of $A$-modules $\widehat{M}\otimes_{A}M\rightarrow A$ given by $f\otimes x\mapsto f(x)$ is an isomorphism where the $A$-module $\widehat{M}=\operatorname{Hom}_{A}(M,A)$ is called the dual of $M$. In this case, $\widehat{M}$ is also a finitely generated projective $A$-module of rank 1.
All these results are standard and can be found in the literature (see e.g. [1], [10], [13] and [16]).
For any commutative ring $A$, by $\operatorname{Pic}(A)$ we mean the set of isomorphism classes of invertible modules over $A$ (invertible modules are also called line bundles). This set $\operatorname{Pic}(A)$ by the operation $L_{1}]\cdot[L_{2}]=[L_{1}\otimes_{A}L_{2}]$ is an Abelian group whose identity element is the isomorphism class of $A$ and the inverse of each element $L]\in\operatorname{Pic}(A)$ is the isomorphism class of $\widehat{L}$. The group $\operatorname{Pic}(A)$ was introduced by Grothendieck and given the name Picard group of $A$. We will often denote the isomorphism class $[L]$ simply by $L$ if there is no confusion.
@@
It can be seen that the Picard group commutes with direct limits. More precisely, if $(A_{k},f_{i,k})$ is a direct system of rings over a (directed) poset $I$, then the groups $\operatorname{Pic}(A_{k})$ together with the maps $\operatorname{Pic}(f_{i,k})$, as transition morphisms, is a direct system of Abelian groups over the same (directed) poset and we have the canonical isomorphism of groups $\operatorname{Pic}(\lim\limits_{\overrightarrow{k\in I}}A_{k})\simeq\lim\limits_{\overrightarrow{k\in I}}\operatorname{Pic}(A_{k})$.
The Picard group also commutes with “finite” direct products of rings. In particular, for given rings $A$ and $B$ consider their direct product ring $R=A\times B$, then the morphism of groups $\operatorname{Pic}(R)\rightarrow\operatorname{Pic}(A)\times\operatorname{Pic}(B)$ given by $L\mapsto(L\otimes_{R}A,L\otimes_{R}B)$ is an isomorphism. In this regard see also Theorem 4.6.
@@
Let $I$ be an ideal of a ring $A$ and consider the canonical ring map $f:A\rightarrow A/I$. If $I$ is contained in the Jacobson radical of $A$, then the map $\operatorname{Pic}(f):\operatorname{Pic}(A)\rightarrow\operatorname{Pic}(A/I)$ is injective. If moreover, $I$ is contained in the nil-radical of $A$, then the map $\operatorname{Pic}(f)$ is an isomorphism. In particular, we have a canonical isomorphism of groups $\operatorname{Pic}(A)\simeq\operatorname{Pic}(A_{\mathrm{red}})$.
Parts of the proof can be found in [1, Chap III, Proposition 2.12] and [5, Chap X, §5, Theorem 5.10, 1.(b)]), but finding an integrated and single proof of this theorem, especially the surjectiveness of the map $\operatorname{Pic}(f)$, is a rather difficult task. Hence, we provide a proof for it.
First recall that if $A\rightarrow B$ is a surjective morphism of rings then for any $B$-modules $M$ and $N$ we have $\operatorname{Hom}_{B}(M,N)=\operatorname{Hom}_{A}(M,N)$.
It is well known that if $M$ and $N$ are finitely generated projective $A$-modules such that $M/IM\simeq N/IN$ and $I$ is contained in the Jacobson radical of $A$, then $M\simeq N$. This establishes the injectivity of $\operatorname{Pic}(f)$.
To see surjectivity, let $L$ be an invertible module over $A/I$. Then there exists a morphism of $A$-modules $g:(A/I)^{n}\rightarrow(A/I)^{n}$ with $n\geqslant 1$ such that $L\simeq\operatorname{Im}(g)$ and $g$ is idempotent, i.e., $g\circ g=g$ (because there is a surjective morphism of $A$-modules $\varphi:(A/I)^{n}\rightarrow L$, and since $L$ is a projective module there is a morphism of $A$-modules $\psi:L\rightarrow(A/I)^{n}$ such that $\varphi\circ\psi$ is the identity map, then take $g:=\psi\circ\varphi$). The induced ring map between the rings of $n\times n$ square matrices $\mathscr{M}_{n}(f):\mathscr{M}_{n}(A)\rightarrow\mathscr{M}_{n}(A/I)$ is surjective and $\operatorname{Ker}(\mathscr{M}_{n}(f))=\mathscr{M}_{n}(I)$, the set of all $n\times n$ matrices with entries in $I$.
We claim that the two-sided ideal $\mathscr{M}_{n}(I)$ is contained in the set of nilpotent elements of $\mathscr{M}_{n}(A)$. Indeed, take a matrix $M$ in $\mathscr{M}_{n}(I)$, and let $J$ be the ideal of $A$ generated by the entries of $M$. Then $A\in\mathscr{M}_{n}(J)$. Clearly $J$ is a nilpotent ideal, i.e., $J^{d}=0$ for some $d\geqslant 1$ (because $J$ is a finitely generated ideal contained in the nilradical of $A$). For any ideals $I_{1}$ and $I_{2}$ of $A$ we have $\mathscr{M}_{n}(I_{1})\mathscr{M}_{n}(I_{2})\subseteq\mathscr{M}_{n}(I_{1}I_{2})$. This shows that $A^{d}=0$. This establishes the claim.
It is well known that if a two-sided ideal of a ring (not necessarily commutative) is contained in the set of nilpotent elements, then the idempotents of this ring can be lifted modulo this ideal. This shows that the natural ring map $\mathscr{M}_{n}(f)$ lifts idempotents.
We know that there is a natural isomorphism of rings $\mathscr{M}_{n}(A)\rightarrow\operatorname{End}_{A}(A^{n})$ which is given by $M\mapsto T_{M}$ where $T_{M}:A^{n}\rightarrow A^{n}$ is defined as $x\mapsto Mx$. Also note that if $M$ is an $A$-module, then any ring map $A\rightarrow B$ induces the ring map $\operatorname{End}_{A}(M)\rightarrow\operatorname{End}_{B}(M\otimes_{A}B)$ that is given by $h\mapsto h\otimes 1_{B}$.
In particular, the following diagram of rings is commutative:
[…]
Hence, the natural ring map $\operatorname{End}(A^{n})\rightarrow\operatorname{End}((A/I)^{n})$ lifts idempotents. Thus there exists a morphism of $A$-modules $h:A^{n}\rightarrow A^{n}$ such that $h\circ h=h$ and $h\otimes 1_{A/I}=g$ with considering the identification $(A/I)^{n}=A^{n}\otimes_{A}A/I$. Since $h$ is idempotent, we have a decomposition of $A$-modules $A^{n}\simeq\operatorname{Ker}(h)\oplus\operatorname{Im}(h)$. This shows that $M:=\operatorname{Im}(h)$ is a finitely generated projective $A$-module. So we have the following split exact sequence of $A$-modules:
${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${K\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${A^{n}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{h^{\prime}}$${M\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0}$
where $h^{\prime}(x):=h(x)$ and $K=\operatorname{Ker}(h^{\prime})$. Then tensoring with $\otimes_{A}A/I$ leaves the sequence split exact:
${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${K\otimes_{A}A/I\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${A^{n}\otimes_{A}A/I\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{h^{\prime}\otimes 1}$${M\otimes_{A}A/I\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0.}$
This shows that $M/IM\simeq\operatorname{Im}(h^{\prime}\otimes 1)$. Note that $h=i\circ h^{\prime}$ where $i:M\rightarrow A^{n}$ is the inclusion map. We have $h\otimes 1=(i\otimes 1)\circ(h^{\prime}\otimes 1)$. The map $i\otimes 1$ is injective because, similar to the above, the split exact sequence:
${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${M\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{i}$${A^{n}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{Im}(h)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0}$
gives the following split exact sequence:
${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${M\otimes_{A}A/I\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{i\otimes 1}$${A^{n}\otimes_{A}A/I\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{Im}(h)\otimes_{A}A/I\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0.}$
Thus $\operatorname{Im}(h\otimes 1)\simeq\operatorname{Im}(h^{\prime}\otimes 1)$. Then $M/IM\simeq\operatorname{Im}(h\otimes 1)\simeq\operatorname{Im}(g)\simeq L$. Finally, we need to show that the rank map of $M$ is the constant function 1. If $\mathfrak{p}$ is a prime ideal of $A$, then $I\subseteq\mathfrak{p}$. Thus $M_{\mathfrak{p}}/I_{\mathfrak{p}}M_{\mathfrak{p}}\simeq L_{\mathfrak{p}/I}\simeq A_{\mathfrak{p}}/IA_{\mathfrak{p}}$ is a cyclic module over $A_{\mathfrak{p}}/IA_{\mathfrak{p}}$. So by the Nakayama lemma, we get that $M_{\mathfrak{p}}$ is a nonzero cyclic $A_{\mathfrak{p}}$-module. But $M_{\mathfrak{p}}$ is also a free $A_{\mathfrak{p}}$-module. This shows that $M_{\mathfrak{p}}\simeq A_{\mathfrak{p}}$.